Circular No. 8379

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004dj IN NGC 2403 C. J. Stockdale, Marquette University; R. A. Sramek, National Ra-

dio Astronomy Observatory; K. W. Weiler, Naval Research Laboratory; S. D. Van Dyk, Infrared Processing and Analysis Center, California Institute of Technology; N. Panagia, European Space Agency and Space Telescope Science Institute; D. Pooley and W. Lewin, Massachusetts Institute of Technology; and J. M. Marcaide, University of Valencia, report the detection of radio emission from the type-IIP supernova 2004dj (IAUC 8377, 8378) with the Very Large Array (VLA) of the National Radio Astronomy Observatory, with a radio-flux density of 1.77 ± 0.13 mJy at 8.46 GHz (wavelength 3.6 cm) and upper limits (3σ) of 1.4 mJy at 22.46 GHz (wavelength 1.3 cm) and 2.9 mJy at 43.34 GHz (wavelength 0.7 cm) on Aug. 2.9 UT in D configuration. The 8.46-GHz detection was measured at the following position: α = 7^{h}37^{m}16^{s}.916, δ = +65^{o}35′56′′.97 (equinox 2000.0; ± 0′′.2 in each coordinate). This is in close agreement with the reported optical position (IAUC 8377). Previous observations of NGC 2403 on 2003 Jan. 31 showed no radio emission near the supernova position, with upper limits (3σ) of 0.41 mJy at 8.4 GHz and 0.31 mJy at 22.5 GHz. VLA monitoring of this source is continuing.

Visual magnitude estimates of SN 2004dj: Aug. 3.111 UT, 11.9 (P. Schmeer, Bischmisheim, Germany); 3.847, 11.7 (K. Hornoch, Lelekovice, Czech Republic); 4.041, 11.7 (M. Lehky, Hradec Kralove, Czech Republic).

SUPERNOVA 2004dk IN NGC 6118 F. Patat and G. Pignata, European Southern Observatory; and S.

Benetti, Istituto Nazionale di Astrofisica, on behalf of the ERTN (see IAUC 8341), report that a fully reduced spectrum (range 320–900 nm; resolution 1.2 nm) of SN 2004dk (cf. IAUC 8377), obtained on Aug. 3.85 UT by J. Aceituno at the Calar Alto Observatory 2.2-m telescope (+ CAFOS), shows it to be a type-Ic supernova. The spectrum is similar to that of SN 1994I (Filippenko et al. 1995, Ap.J. 415, L11) a few days before maximum light, showing transitions from Ca II, Fe II, Na I, Si II, and O I, all accompanied by P-Cyg profiles. The photospheric expansion velocity, deduced from the minimum of the Si II 635.5-nm absorption trough, is ∼ 9100 km/s.

2004 August 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green